OpenStudy (anonymous):

What is the slope of the line passing through the points (6, -4) and (8, -7)?

OpenStudy (usukidoll):

any attempts?

OpenStudy (danjs):

\[slope = \frac{ -7 - (-4) }{ 8-6 }\] The difference in y coordinate, divided by the difference in the x coordinate is the slope of the line.

OpenStudy (anonymous):

well kinda, i never really understood this, even tho i have an 89 in math

OpenStudy (anonymous):

A. -3/2 B. -2/3 C. -11/14 D. 11/2

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

look above and simplify the fraction,

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

-3/2?

OpenStudy (danjs):

right difference in the y coordinates, -7 - (-4) = -7 + 4 = -3 divided by difference in the x coodinates, 8 - 6 = 2

OpenStudy (anonymous):

Thank you !

jagr2713 (jagr2713):

\[\huge \color {Blue}{Welcome \to OpenStudy!}\] @StonerWill

OpenStudy (danjs):

haha nice mod boarder looking pic @jagr2713 , clever you

jagr2713 (jagr2713):

:D thanks